# Cyber Defense Reinvented: Large Language Models as Threat Intelligence Copilots

### v1

@@ Abstract
The exponential growth of cyber threat knowledge, exemplified by the expansion of databases such as MITRE-CVE and NVD, poses significant challenges for cyber threat analysis. Security professionals are increasingly burdened by the sheer volume and complexity of information, creating an urgent need for effective tools to navigate, synthesize, and act on large-scale data to counter evolving threats proactively. However, conventional threat intelligence tools often fail to scale with the dynamic nature of this data and lack the adaptability to support diverse threat intelligence tasks.
@@
In this work, we introduce CyLens, a cyber threat intelligence copilot powered by large language models (LLMs). CyLens is designed to assist security professionals throughout the entire threat management lifecycle, supporting threat attribution, contextualization, detection, correlation, prioritization, and remediation. To ensure domain expertise, CyLens integrates knowledge from 271,570 threat reports into its model parameters and incorporates six specialized NLP modules to enhance reasoning capabilities. Furthermore, CyLens can be customized to meet the unique needs of different organizations, underscoring its adaptability. Through extensive evaluations, we demonstrate that CyLens consistently outperforms industry-leading LLMs and state-of-the-art cybersecurity agents. By detailing its design, development, and evaluation, this work provides a blueprint for leveraging LLMs to address complex, data-intensive cybersecurity challenges.
1 Introduction
@@
We are living in an era of rapid digital transformation, where technological advancements are closely linked to the growing prevalence of cyber threats. In recent years, the landscape of cyber threats has evolved dramatically, marked by a 25% annual increase in reported CVEs [40]. The increase in CVEs can be attributed to several factors, especially the growing complexity of IT systems [76], the widespread adoption of open-source software [22, 79], and the accelerated pace of development cycles [83]. Together, these factors facilitate the exploitation of cyber threats in modern software ecosystems.
To enable efficiently analyze and counteract these threats, Cyber Threat Intelligence (CTI) [23] has emerged as a crucial strategy. CTI provides actionable insights derived from the analysis of threat data and empowers security teams to anticipate, identify, and mitigate cyber risks effectively [89, 110, 8].
@@
However, CTI faces critical challenges in scalability and adaptability. First, the overwhelming volume of threat data creates significant bottlenecks [1]. Security teams are often inefficiently processing and correlating this massive influx of security knowledge, limiting their ability to respond to emerging cyber threats. Second, integrating CTI into existing cybersecurity platforms often requires extensive customization, which poses a significant barrier for small and medium-sized organizations who lack resources to deploy CTI effectively. To address these challenges, it is essential to propose advanced CTI solutions capable of scalably collecting and analyzing extensive cybersecurity knowledge while adaptively support customization tailored to diverse organizational needs.
Figure 1: (a) The lifecycle of threat management. Each task consists of multiple analytical targets, as detailed in Table 1. (b) The development methodology of CyLens, as detailed in section 3 and section 4.
Large language models (LLMs) have recently demonstrated exceptional capabilities, driven by their scalability to massive knowledge through pre-training and their adaptability for organization-specific applications via fine-tuning. Beyond these strengths, LLMs can be further augmented with task-oriented prompts or external modules, enabling more advanced analyses that require topic modeling [96], entity and relation recognition [97, 37], and retrieval-augmented generation (RAG) [33, 10]. Hence, LLMs emerge as a promising solution for assisting humans in specialized fields such as bug fixing [72, 94] and fuzz testing [38, 104, 24]. Consequently, we raise a critical question: how can LLMs be systematically developed as copilots to support CTI activities?
This work. To address this question, we propose CyLens, an LLM-powered copilot system designed to enhance CTI. CyLens consists of several sizes—1B, 8B, and 70B—to accommodate diverse needs and computational constraints. We characterize CyLens by the following aspects:
Lifecycle-Oriented Threat Management – As illustrated in Figure 1-(a), CyLens supports six tasks that cover the entire lifecycle of threat management. To analyze the properties of threats, CyLens enables ❶ attribution by identifying threat actors and gathering technical evidence, such as TTPs and exploitation paths, to trace the origin of a threat. It also facilitates ❷ contextualization by enriching metadata—such as affected systems—to provide a more comprehensive understanding of the threat landscape. Next, ❸ detection involves specifying available identifiers, such as CVE [14], while ❹ correlation enables CyLens to uncover adversarial networks by identifying relationships between the current threat and others. CyLens then drives ❺ prioritization by quantifying the severity and predicting the likelihood of exploitation. Building on this analytical foundation, ❻ remediation effectively counteracts threats and limits their broader impact.
Collective Expertise. As shown in Figure 1-(b), CyLens manages web-scale CTI knowledge through a combination of development methodologies. Our approach is motivated by the fact that cyber threats co-exist as an interdependent community, where attack vectors, infrastructure, and exploitations are deeply interconnected [46, 6, 93, 12, 103]. Effectively addressing these complex relationships requires a vast, integrated set of expertise rather than isolated knowledge.
To achieve this, CyLens first undergoes pre-training on a diverse set of databases—MITRE-CVE [14], CWE [16], CAPEC [17], ATT&CK [18], NVD [59], Exploit-DB [62], and correlated third-party repositories (e.g., [39, 5])—to incorporate extensive CTI knowledge from 271,570 threat reports. We introduce a curriculum pre-training approach, where “curriculum” data is designed to progressively embed CTI expertise into CyLens. Following pre-training, CyLens undergoes fine-tuning using task-oriented datasets addressing critical CTI tasks (❶-❻). To enhance its analytical depth, we employ a cascading reasoning process, allowing CyLens to systematically build awareness of dependencies within CTI analysis. Additionally, we integrate specialized design modules to reinforce CyLens ’s ability to conduct rigorous reasoning, escalating it as a human-usable CTI copilot.
@@
Operational Proficiency. Our experiments demonstrate CyLens’s effectiveness across all above tasks under two categories of cyber threats: (i) Historical Threats—previously recorded threats that security analysts may not be fully aware of, and (ii) Zero-Day Threats—newly emerging threats that have not yet been documented. Across both threat categories, CyLens consistently outperforms SOTA models and industry-leading LLMs (e.g., ChatGPT and Gemini), underscoring its versatility in handling diverse cybersecurity challenges while excelling in proactive threat hunting within the ever-evolving cyberattack landscape.
Adaptive to Customization. CyLens also demonstrates adaptability in addressing organization-specific threats related to particular products, vendors, or threat groups. As shown in our evaluations in section 5.6, CyLens consistently delivers superior and balanced performance across diverse organizational requirements. This adaptability is particularly valuable when an organization’s threat-related data volume is insufficient, enabling effective CTI analysis even in data-scarce scenarios.
Contributions. We summarize following contributions:
    • LLM-as-CTI-Copilot: We propose CyLens, a scalable and adaptive system for CTI practice. CyLens is designed to accommodate diverse needs, making it well-suited for both research and industry applications.
@@
    • Empirical Insights: Through extensive evaluations, we not only demonstrate CyLens’s effectiveness, but also reveal the varying performance of SOTA and industry-leading LLMs, highlighting the inherent complexity of CTI analysis and difficulty of achieving consistent expertise.
@@
    • Blueprint Perspective: The design and development of CyLens serve as a blueprint for integrating LLM copilots into cybersecurity practice. This endeavor not only advances the capabilities of CTI systems but also provides insights that could drive the development of expert-level AI across other high-stakes domains [^1] Codes and datasets are organized at https://anonymous.4open.science/r/security-agent-4BB7/.
2 Background and Related Work
2.1 Cyber Threat Ingelligence (CTI)
@@
AI-for-Security has been a long-standing research area, with extensive studies covering a broad range of cybersecurity topics, including anomaly detection [109, 68], malware classification[47, 69], and intrusion detection[48, 31]. Among these, CTI has emerged as a systematic approach to threat management, enabling organizations to collect, analyze, and operationalize intelligence on cyber threats [20, 95, 89].
@@
Although there is no universal specification for CTI practices, we categorize CTI into six critical threat management tasks based on extensive literature [82, 32, 86, 54, 42, 75, 89, 71, 114, 43] and the MITRE initiative [3], a foundational effort in the field. As illustrated in Figure 1-(a), the lifecycle of threat intelligence begins with the extraction and analysis of threat properties (❶ & ❷) [42, 105, 75], progresses to the identification of adversary community (❸ & ❹) [32, 86, 54, 114], and culminates in the development of effective countermeasures (❺ & ❻) [89, 71, 43]. In this work, we set the scope of the CTI tasks as follows:
@@
    • Attribution–Identifying the origin of cyber threats by profiling threat actors, analyzing tactics, techniques, and procedures (TTPs), and tracing exploitation paths to link cyber incidents to specific adversaries or groups. Attribution helps organizations understand attacker intent, capabilities, and historical attack patterns [105, 42].
@@
    • Contextualization–Enriching raw threat intelligence with relevant metadata, such as affected systems, attack infrastructures, and potential impacts. Contextualization enables security analysts to assess threats with a comprehensive picture of the threat landscape[52, 75].
    • Detection–Identifying cyber threats using available indicators, such as CVE [18] and CWE [16]. Detection assists the recognition of both known and emerging threats [114, 32].
    • Correlation–Connecting disparate threats to uncover hidden relationships between attack campaigns and disclose the broader landscape of adversarial community. Correlation enhances situational awareness and strengthens incident investigations by providing a more comprehensive view of threat activities [86, 54, 52].
@@
    • Prioritization– Ranking threats based on severity, exploitability, and potential impact to determine the urgency of remediation. Effective prioritization optimizes resource allocation by focusing on the most critical threats [89, 43].
@@
    • Remediation–Developing and recommending actionable countermeasures, including security patches, mitigation strategies, and responsible advisories. Remediation ensures that organizations can effectively neutralize threats and reinforce their security posture[43, 74].
@@
Within each CTI task, we define distinct analytical targets. For instance, prioritization involves assessing both severity, measured by the Common Vulnerability Scoring System (CVSS) [60], and exploitation potential, measured by the Exploit Prediction Scoring System (EPSS) [29]. The detailed analytical targets for each CTI task are presented in Table 1 and are evaluated extensively in section 5.
2.2 LLM-as-Copilot
Figure 2: Development pipeline of LLM copilots.
Due to their vast inherent knowledge and ability to handle large-scale downstream tasks, LLMs have been extensively adapted for domain-specific applications, an effort referred to as “LLM-as-Copilot” [51, 34, 90]. Unlike the broader concept of an "agent," which is designed to perform human-like behaviors, a copilot specifically emphasizes its role as a human assistant [85, 55, 88]. This distinction is particularly critical in high-stakes domains such as cybersecurity [112, 28], healthcare [4, 98, 108], and finance [107, 111, 70], where expert guidance and decision support are essential.
Training Methodology. As illustrated in Figure 2, the development of LLM copilots follows a pre-training and fine-tuning trajectory. Pre-training leverages large-scale corpora to embed extensive domain expertise into base LLMs through a causal language modeling objective [9, 25, 45, 27]. Subsequently, fine-tuning incorporates environment-aware instructions and utilizes supervised learning [61, 78] to refine the model’s reasoning capabilities for practical applications.
This pre-training followed by fine-tuning approach has proven particularly effective in agent development requiring broad domain expertise, such as existing efforts in AI-assisted programming [2, 91]. Similarly, the interconnected nature of cyber threats necessitates a deep and expansive CTI knowledge. Thus, this work adopts the same development pipeline to enhance LLM capabilities, advancing them toward the creation of an effective and practical cybersecurity copilot.
Prompting. Besides training methodologies, inference-time prompting plays a crucial role in regulating an LLM’s behavior during execution [63, 36]. While developing LLM copilots cannot rely solely on prompting, it provides flexibility by enabling them to interact with real-world resources, such as functional tools [87] or retrieving web-based knowledge [53].
In this work, we enhance CyLens’s inference by implementing 6 NLP modules, integrated through either additional functions or strategic prompting techniques. These enhancements improve CyLens’s reasoning, contextual awareness, and overall effectiveness in CTI analysis.
3 Data Management and Pre-Training
In this section, we describe the collection and preprocessing of threat reports used in pre-training (section 3.1) and introduce a curriculum-based pre-training approach that integrates large-scale cyber threat knowledge into CyLens (section 3.2). The subsequent section (section 4) details the fine-tuning and inference design.
3.1 Threat Data: Collection and Preprocessing

### v2

# CyLens: Towards Reinventing Cyber Threat Intelligence in the Paradigm of Agentic Large Language Models

@@ Abstract.
The exponential growth of cyber threat knowledge, exemplified by the expansion of databases such as MITRE-CVE and NVD, poses significant challenges for cyber threat analysis. Security professionals are increasingly burdened by the sheer volume and complexity of information, creating an urgent need for effective tools to synthesize massive knowledge and counter evolving threats proactively. However, conventional threat intelligence often struggle to scale with the rapidly expanding threat landscape and lack the adaptability required to support diverse threat intelligence tasks.
@@
In this work, we introduce CyLens, a cyber threat intelligence system powered by large language models (LLMs). CyLens is designed to assist security professionals throughout the entire threat analysis lifecycle, supporting threat attribution, contextualization, correlation, prioritization, and remediation. To ensure domain expertise, CyLens integrates knowledge from 271,570 threat reports into its model parameters and incorporates six specialized NLP modules to enhance reasoning capabilities. Furthermore, CyLens can be customized to meet the unique needs of different organizations, underscoring its adaptability. Through extensive evaluations, we demonstrate that CyLens consistently outperforms industry-leading LLMs and state-of-the-art cybersecurity agents. By detailing its design, development, and evaluation, this work provides a blueprint for leveraging LLMs to address complex, data-intensive cybersecurity challenges.
Cyber Threat Intelligence, Large Language Models, Threat Hunting
* Xiaoqun Liu and Jiacheng Liang contributed equally to this work.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Systems security
1. Introduction
@@
We are living in an era of rapid digital transformation, where technological advancements are closely linked to the growing prevalence of cyber threats. In recent years, the landscape of cyber threats has evolved dramatically, marked by a 25% annual increase in reported Common Vulnerabilities and Exposures (CVEs) (Intelligence, 2023). The increase in CVEs can be attributed to several factors, especially the growing complexity of IT systems (Qualys, 2024), the widespread adoption of open-source software (Dam and Neumaier, 2023; Reading, 2024), and the accelerated pace of development cycles (Security, 2024). Together, these factors facilitate the exploitation of cyber threats in modern software ecosystems.
Figure 1. (Left) An agentic LLM execution on threat intelligence tasks. (Right) The customization workflow to enhance the CTI proficiency of pre-trained LLMs.
[…]
Introduction Example.
@@
To efficiently analyze and counteract these threats, Cyber Threat Intelligence (CTI) (Dehghantanha et al., 2018) has emerged as a crucial strategy. CTI provides actionable insights derived from the analysis of threat data and empowers security teams to anticipate, identify, and mitigate cyber risks effectively (Sun et al., 2023; Bromiley, 2016). However, it faces two critical challenges: (1) Scalability: The overwhelming volume of threat data creates significant bottlenecks (Abu et al., 2018). Security teams are often inefficiently processing and correlating this massive influx of security knowledge, limiting their ability to respond to emerging cyber threats. (2) Adaptability: Integrating CTI into existing cybersecurity platforms often requires extensive customization, which poses a significant barrier for small and medium-sized organizations who lack resources to deploy CTI effectively. To address these challenges, it is essential to propose advanced CTI solutions that adaptively support customization tailored to diverse organizational needs.
Figure 2. The lifecycle of threat intelligence tasks and running examples.
[…]
CTI pipeline and example.
Large language models (LLMs) have recently demonstrated exceptional capabilities in natural language understanding and reasoning. Compared to conventional CTI solutions, such as rule-based systems (Ali and Peng, 2024; Sun et al., 2023) or function-frozen AI models (Montasari et al., 2021; Kumar et al., 2024; Kant and Amrita, 2023; Dutta and Kant, 2020), LLMs offer superior scalability, as they can generalize across diverse knowledge-intensive domains without the need for task-specific architectures or manually crafted features (Budnikov et al., 2024; Kunter et al., 2013; Brown, 2020). LLMs also exhibit strong adaptability. Techniques like instruction tuning and few-shot learning enable them to respond effectively to dynamic and data-scarce environments, where traditional models often struggle to keep pace (Liu et al., 2024; Paeedeh et al., 2024; Wang et al., 2021; Al-Shedivat et al., 2017).
However, despite recent studies showing that industry-leading LLMs are not yet fully qualified for CTI analysis (Srikanth et al., 2024; Kouremetis et al., 2025; Xu et al., 2024; Shafee et al., 2024), the blueprint for systematically customizing LLMs to support scalable and adaptive CTI analysis remains largely unexplored. This leads us to a central question: How can LLMs be systematically developed to support diverse CTI activities?
Our work. To address this question, we propose CyLens, an LLM-powered agentic system designed to enhance CTI. CyLens is available in multiple model sizes—1B, 8B, and 70B—to accommodate varying computational constraints. As illustrated in Figure 1, CyLens follows the Agentic LLM paradigm, inheriting pre-trained language understanding and reasoning capabilities, and undergoing a series of targeted customizations to improve domain-specific expertise. We characterize CyLens by the following aspects:
Lifecycle-Oriented Threat Analysis – CTI practice encompasses a lifecycle of analyzing and counteracting cyber threats (Irshad and Siddiqui, 2023; Xiang et al., 2024; Qiang et al., 2017; Sun et al., 2023; Patil and Malla, 2022). As illustrated in Figure 2, CyLens is aligned with this lifecycle and supports five groups of analytical tasks. To investigate the nature and origin of threats, CyLens enables ❶ attribution by identifying threat actors and collecting technical evidence, such as TTPs and campaign, to trace the source of an attack. It further facilitates ❷ contextualization by enriching metadata, including affected systems and infrastructures, to provide a comprehensive understanding of the threat landscape. Next, ❸ correlation involves linking known identifiers, such as CVEs (Corporation, nd), while also uncovering adversarial networks by identifying related or concurrent threats. CyLens supports ❹ prioritization by assessing the severity of threats and estimating the likelihood of exploitation, enabling organizations to allocate resources efficiently. Finally, based on these analytical foundations, ❺ remediation supports the development of actionable countermeasures to effectively neutralize threats and reduce their broader impact.
Scalable Knowledge InfusionCyLens incorporates extensive CTI knowledge from a diverse set of threat databases and repositories, including MITRE-CVE (Corporation, nd), CWE (Corporation, 2006), CAPEC (Corporation, 2007), ATT&CK (Corporation, 2013), NVD (National Institute of Standards and Technology, 2025), Exploit-DB (Offensive Security, 2025), and their correlated threat reporting platforms (e.g., IBM-XForce (IBM X-Force, 2025), Apache (Apache Software Foundation, 2025a), and Red Hat (Red Hat, Inc., 2025)), resulting in a collection of 271,570 threat reports.
To integrate extensive CTI knowledge, CyLens adopts a curriculum training methodology, where the “curriculum” consists of ordered threat reports designed to progressively embed CTI expertise into the model. This approach enables scalable knowledge infusion by allowing the training process to incorporate both extensive historical data and emerging threats. By adjusting pacing steps and balancing newer data with earlier content, curriculum training eliminates the need for retraining from scratch.
Adapting through TuningCyLens exhibits strong adaptability in addressing specific analytical needs, such as threat analysis related to particular products, vendors, or threat groups. To enable this, we apply instructional fine-tuning (or simply instruction tuning), which leverages task-oriented datasets aligned with critical CTI tasks (❶–❺). Within these datasets, we introduce a cascading reasoning framework that reflects the inherent task dependencies across the CTI lifecycle. This structured design allows CyLens to emulate best practices as expert-level CTI analysis.
Modularized Inference – At inference time, we integrate specialized design modules to enhance CyLens’s capacity for rigorous and reliable reasoning. For example, threat actor attribution involves applying named entity recognition (NER) to threat reports and retrieving relevant, up-to-date metadata regarding recent exploitation activities. By systematically organizing CyLens ’s inference process through these modules, we improve the consistency and precision of its analytical outputs.
Validated Proficiency – In experiments, we collect real-world threat reports from platforms including Oracle Security Alerts (Oracle Corporation, 2025), Red Hat Bugzilla (Red Hat, Inc., 2025), Adobe Security Bulletin (Adobe Systems Incorporated, 2025), and the Apache Mailing List (Apache Software Foundation, 2025a). We design two groups of experiments to evaluate the scalability and adaptability of CyLens. For scalability evaluation, we evaluate CyLens’s effectiveness across a large volume of threats spanning CTI tasks ❶–❺, considering multiple analytical targets under each task. For example, threat contextualization includes identifying affected systems, attack infrastructure, and potential impacts. For adaptability evaluation, we focus on specialized threats targeting specific products (e.g., Acrobat), vendors (e.g., Adobe), or threat categories (e.g., DDoS). These scenarios reflect the diverse and evolving needs of industrial organizations and research institutions (Saeed et al., 2023; Van Haastrecht et al., 2021; Sakellariou et al., 2022).
@@
In both evaluations, we categorize threats into two groups: (i) Historical Threats—previously recorded threats that security analysts may not be fully aware of; and (ii) Zero-Day Threats—newly emerging threats that have not yet been documented. Across both categories, CyLens consistently outperforms state-of-the-art (SOTA) models and industry-leading LLMs (e.g., ChatGPT and Gemini), demonstrating its versatility in addressing a broad range of cybersecurity challenges and its effectiveness in supporting proactive threat hunting within a dynamic threat landscape.
Contributions. We summarize the following contributions:
    • We propose CyLens, a scalable and adaptive agentic system for CTI. CyLens supports lifecycle-oriented threat analysis, encompassing tasks from attribution to remediation. To accommodate varying computational constraints, CyLens is implemented in three model sizes: 1B, 8B, and 70B. Through comprehensive knowledge infusion, followed by instruction tuning and modularized inference, CyLens is equipped with deep CTI expertise and consistently delivers rigorous, reasoning-driven analysis.
@@
    • Through extensive evaluations on historical and zero-day threats, we not only demonstrate CyLens’s effectiveness, but also reveal the varying performance of SOTA and industry-leading LLMs, highlighting the inherent complexity of CTI analysis and difficulty of achieving consistent expertise.
@@
    • The design and development of CyLens serve as a blueprint for integrating agentic LLM into cybersecurity practice. This endeavor not only advances the capabilities of CTI systems but also provides insights that could drive the development of expert-level AI across other high-stakes domains. [^1] Codes and datasets are organized at https://anonymous.4open.science/r/security-agent-4BB7/
2. Background and Related Work
2.1. Cyber Threat Ingelligence (CTI)
@@
AI-for-Security has been a long-standing research area, with extensive studies covering a broad range of cybersecurity topics, including anomaly detection (Yaseen, 2023; Pang et al., 2021), malware classification(Kalash et al., 2018; Pascanu et al., 2015), and intrusion detection(Khraisat et al., 2019; Frank, 1994). Among these, CTI has emerged as a systematic approach in cybersecurity, enabling organizations to collect, analyze, and operationalize cyber threats (Cybersecurity and , CISA; Wagner et al., 2019; Sun et al., 2023).
@@
Although there is no universal taxonomy in CTI, we categorize its involved tasks into five critical aspects based on extensive literature (Gao et al., 2020; Lin et al., 2023; Irshad and Siddiqui, 2023; Qiang et al., 2017; Sun et al., 2023; Patil and Malla, 2022; Zhu et al., 2024b; Jacobs et al., 2020; Kouremetis et al., 2025) and the MITRE initiative (Al-Sada et al., 2024), a foundational effort in the field. As illustrated in Figure 2, the lifecycle of threat intelligence begins with investigatig threat properties (❶ & ❷) (Irshad and Siddiqui, 2023; Xiang et al., 2024; Qiang et al., 2017), progresses to identify adversary community (❸) (Gao et al., 2020; Lin et al., 2023; Zhu et al., 2024b), and culminates in developing countermeasures (❹ & ❺) (Sun et al., 2023; Patil and Malla, 2022; Jacobs et al., 2020). Formally, this work set the scope of the CTI tasks as follows:
@@
    • Attribution–Identifying the origin of cyber threats by profiling threat actors, analyzing tactics, techniques, and procedures (TTPs), and tracing attack campaign to link cyber incidents to specific adversaries or groups. Attribution helps organizations understand attacker intent, capabilities, and historical attack patterns (Xiang et al., 2024; Irshad and Siddiqui, 2023).
@@
    • Contextualization–Enriching relevant metadata, such as affected systems, attack infrastructures, and potential impacts. Contextualization enables security analysts to assess threats with a comprehensive picture of the threat landscape(Li et al., 2022; Qiang et al., 2017).
    • Correlation–Linking available threat identifiers, such as CVE (Corporation, 2013) and CWE (Corporation, 2006), and uncovering related or concurrent activities within the same attack campaigns. Correlation supports to identify both known and emerging threats (Zhu et al., 2024b; Gao et al., 2020), while exposing the broader landscape of the attack community (Lin et al., 2023).
@@
    • Prioritization– Ranking threats based on severity and likelihood of exploitation to determine their remediation urgency. Effective prioritization optimizes resource allocation by focusing on the most critical and high-risk threats.(Sun et al., 2023; Jacobs et al., 2020).
@@
    • Remediation–Developing and recommending actionable countermeasures, including security patches, mitigation strategies, and responsible advisories. Remediation ensures that organizations can effectively neutralize threats and reinforce their security posture (Jacobs et al., 2020; Poehlmann et al., 2021).
@@
Within each CTI task, we define distinct analytical targets. For instance, prioritization involves assessing both severity, measured by the Common Vulnerability Scoring System (CVSS) (National Institute of Standards and Technology (2025), NIST), and exploitation potential, measured by the Exploit Prediction Scoring System (EPSS) (FIRST.org, 2025). The detailed analytical targets for each CTI task are presented in Table 2 and are evaluated extensively in section 5.
2.2. Agentic LLM
As extensively studied in recent works, the agentic LLM paradigm follows a two-stage development process: $\theta_{\texttt{pt}}\rightarrow\theta_{\texttt{agent}}$ (Xu et al., 2023; Xi et al., 2023; Yang et al., 2024), where a pre-trained LLM $\theta_{\texttt{pt}}$ serves as the foundation due to its inherent language comprehension and reasoning abilities. The model is then customized into $\theta_{\texttt{agent}}$ to enhance its domain-specific expertise. This customization may involve re-training to infuse large-scale domain knowledge into LLMs—such as in programming synthesis tasks (Xu et al., 2023)—or instruction tuning to adapt LLMs to highly specialized applications, such as cardiological diagnostics (Zhou et al., 2024).
During inference, agentic LLMs can be further enhanced using specialized prompts to structure to regularize their reasoning processes (Wei et al., 2022; Hu et al., 2024), or by integrating tool functions (e.g., APIs) to augment their problem-solving capabilities in domain-specific tasks (Qin et al., 2023a, b). For example, given a threat report $d$, one can implement a named entity recognition function $\texttt{NER}(\cdot)$ callable by the LLM, such that $\texttt{NER}(d)\rightarrow\mathcal{E}$ extracts threat-related entities (evidence) $\mathcal{E}$ from $d$. These tool functions may be implemented as embodied interfaces triggered through prompt instructions (Schick et al., 2023; Li et al., 2024).
3. Methodology
3.1. Overview
This section details the design of CyLens. Following the agentic LLM paradigm (section 2.2), CyLens requires both customization and tailored inference to address heterogeneous CTI tasks.
Given the diverse analytical requirements and task-specific disparities in CTI, we adopt a two-stage customization consisting of re-training and instruction tuning to enhance LLMs with CTI expertise. In the re-training stage, we build a large-scale threat corpus (section 3.2) and develop a curriculum training approach (section 3.3) to progressively infuse CTI knowledge into LLMs. Next, we apply instruction tuning using a dataset curated with a cascading reasoning layout (section 3.4), which reflects the dependencies among CTI tasks. A fine-tuned LLM can then be guided to follow this reasoning workflow, enabling it to perform multi-step threat analyses in a structured and expert-like manner.
During inference, CyLens uses six NLP modules to structure its generation workflow (section 3.5). These include topic modeling to determine the query intent, entity and relation extraction, information retrieval, and finally, reasoning and evidence summarization to generate coherent, task-specific responses. Together, these modules help CyLens produce accurate, structured outputs while minimizing hallucinations during open-minded generation.
3.2. Threat Corups Construction